Suppose the price of Doritos fall from R12.50 to R8.50. As a result, the quantity of peanuts increases from 50 to 200. Calculate the cross-price elasticity using the arc formulae and interpret your answer.
1
Expert's answer
2019-10-07T10:03:24-0400

Cross-price elasticity is calculated as E=((200-50)/50)/((8.5-12.5)/12.5)=3/(-0.32)=-9.375

Itr means that the quantity of peanuts increses by 9.375% when the price of Doritos decreases by 1%.

The goods are complements, because the coefficient is negative.